OpenStudy (janu16):

You are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x) = –x4 + x3 + 7x2 − x − 6. Describe to the CEO what the graph looks like and, in general, how to sketch the graph without using technology. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0).

OpenStudy (janu16):

@zepdrix

zepdrix (zepdrix):

\[\large\rm P(x)=-x^4+x^3+7x^2-x-6\]

zepdrix (zepdrix):

Do you remember the basic shape of your polynomials? The leading term (largest degree of x) tells us the `degree of polynomial`. If the degree is `even` (which it is in this case), then the "legs" of our function will go off in the same direction, like in the graph of x^2. For `odd` degree, the "legs" go off in opposite directions like in the graph of x^3.

zepdrix (zepdrix):

By "legs" I mean, the end behavior of the function. So we expect the function to go UP at both ends, or DOWN at both ends. Which do you think it will be? Our leading term is `negative`. That should give a hint.

OpenStudy (janu16):

yea

OpenStudy (janu16):

down i think

zepdrix (zepdrix):

Ok good. We don't know exactly what the function looks like just yet, but we know what's happening on the ends of it.

OpenStudy (janu16):

yess

zepdrix (zepdrix):

\[\large\rm P(x)=-x^4+x^3+7x^2-x-6\]They want us to find the break-even point, when the profit function is 0,\[\large\rm 0=-x^4+x^3+7x^2-x-6\]This means we're looking for the x values that make this true. These values go by many names: x-intercepts, roots, zeroes, etc..

zepdrix (zepdrix):

So umm, hmm maybe we can factor this thing :d

OpenStudy (janu16):

yes so weset set p(x0 to 0 and find x by factoirng

zepdrix (zepdrix):

Hmm it's a 4th degree, I'm not sure how easy it would be to factor it. We might want to instead, find a nice simple root to start with, and then apply long division.

zepdrix (zepdrix):

have you done `polynomial long division` or `synthetic division`?

OpenStudy (janu16):

ya i have done synthetic division

zepdrix (zepdrix):

lol you were right the first time on the spelling XD

OpenStudy (janu16):

oh lol

zepdrix (zepdrix):

To do division, we'll first need something that we can divide out. Plug x=1 or x=-1 into the function. See if either of those works out to be true.

OpenStudy (janu16):

what do you suggest? x=1? or x= -1?

zepdrix (zepdrix):

x=1 is always a good place to start :) nice and easy. if it doesn't work, proceed to x=-1, if that doesn't work, we can try something else.

OpenStudy (janu16):

i tink 1 will work so we divide eveyrthing by 1?